A certain number is subtracted 4 times from twice from square of the number and the result is 70. Find the number
Answer by ikleyn(53009)

.
A certain number is subtracted 4 times from twice from square of the number
and the result is 70. Find the number.
~~~~~~~~~~~~~~~~~~


            I edited your post in order for to make sense from nonsense.
            The edited problem is  THIS:

    +-----------------------------------------------------------------+
    |   A certain number is subtracted 4 times from twice the square  |
    |      of the number and the result is 70. Find the number.       |
    +-----------------------------------------------------------------+


                         Solution


Let n be the (or "a") number under the problem's question.

Then, according to the condition, we have this equation

    2n^2 - 4n = 70,

    n^2 - 2n - 35 = 0,

    (n-7)*(n+5) = 0.


So, there are two roots and two possible answers: the number is either -5 or 7.

Solved.